the length of a rectangle is 24 cm and the width is 10cm. how long is the diagonal?
Answer by ewatrrr(24785)

Pythagorean Theorem
.
24cm^2 + 10cm^2 = D^2
676cm^2 = D^2 (tossing out negative solution for unit measure)
26cm = D